A pharmacist has a 40% acid solution and 25% acid solution. How many liters of each must be mixed to form 85 liters of a 31% acid solution?
Answer by stanbon(75887)

A pharmacist has a 40% acid solution and 25% acid solution. How many liters of each must be mixed to form 85 liters of a 31% acid solution?
-------
Equation:
acid + acid = acid
0.40x + 0.25*(85-x) = 0.31*85
----
40x + 25*85 - 25x = 31*85
15x = 6*85
x = (2/5)85
x = 2*17 = 34 liters (amt. of 40% solution needed)
85-x = 51 liters (amt. of 25% solution needed)
================
Cheers,
Stan H.
==================